OpenStudy (anonymous):

Evaluate 9 P 4. 15,120 36 3,024 504 Evaluate 6 C 2. 30 720 120 15

Nnesha (nnesha):

formula for permutation is \[\huge\rm n P_r = \frac{ n! }{ (n-r)! }\] n = 9 r =4 and for combination \[\huge\rm nC_r = \frac{ n! }{ r!(n-r)! }\]